OpenStudy (anonymous):

Find all values of h for which the quadratic equation has no real solutions. =+8x2+3xh0 Write your answer as an equality or inequality in terms of h.

OpenStudy (anonymous):

\[8x^2+3x+h=0\]

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@camerondoherty

OpenStudy (anonymous):

YES H

OpenStudy (anonymous):

HELP ME

OpenStudy (anonymous):

@Abhilash11 help

OpenStudy (mrnood):

from the standard form y = ax^2 + bx + c the 'discriminant' is defined as b^-4ac IF the discriminant is positive or 0 then there are real roots If the discriminant is negative then there are no real roots (You canb see this from the quadratic formula because you cannot have a real sqrt of a negative number)

OpenStudy (anonymous):

Exactly^ this is how you gotta proceed @mathgirl2012

OpenStudy (anonymous):

ok so like the part inside of the square root in the quadratic formula

OpenStudy (anonymous):

See, this is how you'll solve. First pick the values a = B= C =

OpenStudy (mrnood):

\[x= \frac{ -b \pm \sqrt{b ^{2}-4ac} }{2a }\] if b^2-4ac is negative then there is no real answer as you can nsee

OpenStudy (anonymous):

Yes. yes. perfect!

OpenStudy (anonymous):

so yeah ok got it let me try that

OpenStudy (anonymous):

good (Y) :-)

OpenStudy (mrnood):

oops - I missed out a number above ^^ the discriminant is b^2 - 4 ac

OpenStudy (anonymous):

so -23h

OpenStudy (anonymous):

@Abhilash11 @MrNood

OpenStudy (anonymous):

@Abhilash11 @MrNood

OpenStudy (anonymous):

@Abhilash11 @MrNood

OpenStudy (camerondoherty):

Wait, so what exactly is the question?

OpenStudy (anonymous):

finding the h that will make the problem have no solution. The problem is 8x2+3x+h=0

OpenStudy (camerondoherty):

Sorry, I have no idea how to do this cx

OpenStudy (anonymous):

@Abhilash11 @MrNood

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

pls help me @hartnn

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@TheSmartOne what do you think

OpenStudy (anonymous):

@TheSmartOne I will give u a medal and write a testimony it=f you help me

OpenStudy (anonymous):

please @TheSmartOne

TheSmartOne (thesmartone):

So we have... \(\ 8x^2+3x+h=0\)

TheSmartOne (thesmartone):

And sorry that I didn't come right at once. I tend to help like 2-3 people at the same time.. Have to switch tabs...

TheSmartOne (thesmartone):

ok for there to be no real solutions the discriminant has to be less than 0

TheSmartOne (thesmartone):

The discriminant is \(\Large \sqrt{b^2-4ac}\)

TheSmartOne (thesmartone):

And a,b, and c is in the equation \(\Large Ax^2+bx+c=0\) And our equation is \(\Large 8x^2+3x+h\)

TheSmartOne (thesmartone):

So our values for A, B, and C are \(\ A=8\\ B=3\\ C=h\)

TheSmartOne (thesmartone):

So far so good? @mathgirl2012 @~

TheSmartOne (thesmartone):

well just plug in those values into the discriminant and see what value of h would make it less than 0